Which answer shows the next correct step for solving the quadratic equation by taking the square root of both sides?

g^2 - 8g + 16 = 2
(g - 4)^2 = 2
√(g - 4)^2 = √2
g - 4 = √2
|g - 4| = √2
|g - 4|= ±2
±g - 4 = √2

Answer by Alan3354(69443)

Which answer shows the next correct step for solving the quadratic equation by taking the square root of both sides?

g^2 - 8g + 16 = 2
(g - 4)^2 = 2
√(g - 4)^2 = √2
g - 4 = ±√2
g = 4 ±√2
=====================
|g - 4|= ±2
±g - 4 = √2